# Optimal low-rank approximations for linear Gaussian inverse problems on Hilbert spaces Part I: posterior covariance approximation

### v3

@@ B.2 Proofs of Section 4
Given that $xf^{\prime}(x)>0$ for $x\neq 0$, it follows that $f^{\prime}(x)<0$ for $x<0$ and $f^{\prime}(x)>0$ for $x>0$. This implies, by continuity of $f$, that $f^{\prime}(0)=0$. Hence the $f$ has a global minimum only at $x=0$ and $f\geq 0$. Thus, also $\mathcal{L}_{f}\geq 0$. By the Lipschitz assumption on $f^{\prime}$ at 0, there exists $\varepsilon\in(0,1)$ and $M_{0}>0$ such that $f^{\prime}(x)=f^{\prime}(x)-f^{\prime}(0)\leq M_{0}\lvert x\rvert$ for $\lvert x\rvert\leq\varepsilon$. For $\lvert y\rvert\leq\varepsilon$,
@@
$\mathcal{C}^{-1}=\mathcal{C}_{1}^{-1/2}\mathfrak{X}^{-1}\mathcal{C}_{1}^{-1/2}=\mathcal{C}_{1}^{-1/2}\left(I-\sum_{i=1}^{\operatorname{rank}\left({K}\right)}\frac{d_{i}^{2}}{1-d_{i}^{2}}e_{i}\otimes e_{i}\right)\mathcal{C}_{1}^{-1/2}=\mathcal{C}_{1}^{-1/2}(I+ZZ^{*})\mathcal{C}_{1}^{-1/2},$
@@
Let $U,V\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$, $V\not=0$. We have by (19),
$g(U+V)-g(U)=\mathcal{C}_{\textup{pos}}^{1/2}((U+V)(U+V)^{*}-UU^{*})\mathcal{C}_{\textup{pos}}^{1/2}=\mathcal{C}_{\textup{pos}}^{1/2}(VV^{*}+UV^{*}+VU^{*}).$
From $\lVert VV^{*}\rVert=\lVert V\rVert^{2}$, c.f. Lemma A.2, it follows that
$\lVert g(U+V)-g(U)-\mathcal{C}_{\textup{pos}}^{1/2}(UV^{*}+VU^{*})\mathcal{C}_{\textup{pos}}^{1/2}\rVert=\lVert\mathcal{C}_{\textup{pos}}^{1/2}(VV^{*})\mathcal{C}_{\textup{pos}}^{1/2}\rVert\leq\lVert\mathcal{C}_{\textup{pos}}^{1/2}\rVert^{2}\lVert VV^{*}\rVert=\lVert\mathcal{C}_{\textup{pos}}^{1/2}\rVert^{2}\lVert V\rVert^{2}.$
@@
where for second equation we use (29). By definition of $P_{m}$, $\left(P_{m}F^{\prime}(\Lambda(X))\right)_{i}=F^{\prime}(\Lambda(X))_{i}$ for $i\leq m$. Hence,
@@
Let $U,V\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$. Define $\mathcal{Z}\coloneqq\operatorname{ran}{g(U+V)}$ and $\mathcal{W}\coloneqq\{X\in L_{2}(\mathcal{H})_{\mathbb{R}}:\ \operatorname{ran}{X}\subset\mathcal{Z}\}\subset L_{2}(\mathcal{H})_{\mathbb{R}}$. Then $\dim{\mathcal{Z}}<\infty$ by Lemma 4.10 and $\operatorname{ran}{g(U+tV)}\subset\mathcal{Z}$ for all $t\in\mathbb{R}$ by definition (19) of $g$, hence $g(U+tV)\in\mathcal{W}$ for all $t\in\mathbb{R}$. Thus, $F_{f}\circ\Lambda\circ g(U+tV)={\left.\kern-1.2pt(F_{f}\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}\circ g(U+tV)$ for all $t\in\mathbb{R}$. By Lemma 4.11 and Proposition 4.16, ${\left.\kern-1.2pt(F_{f}\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}$ is Fréchet differentiable. By Lemma 4.11, $g$ is Fréchet differentiable. In particular, $g$ is Gateaux differentiable in $U$ in direction $V$. Hence, by the chain rule, $J_{f}$ is Gateaux differentiable at $U$ in the direction $V$. To compute the derivative, we recall that $(e_{j}\otimes e_{k})_{j,k}$ is an ONB of $L_{2}(\mathcal{H})$. The Gateaux derivative of $J_{f}$ at $U$ in the direction $V$ is
@@
Let us write $m\coloneqq\dim{\mathcal{V}}$. For each $n$, let us diagonalise $U_{n}U_{n}^{*}=\sum_{j=1}^{m}\beta_{n,j}\psi_{n,j}\otimes\psi_{n,j}$, where $(\psi_{n,j})_{j=1}^{m}$ forms an ONB of $\mathcal{V}$ and $\beta_{n,j}\geq 0$. Define $j_{n}\coloneqq\operatorname*{arg\,max}_{j\leq m}\beta_{n,j}$, hence $\beta_{n,j_{n}}$ is the largest eigenvalue of $U_{n}U_{n}^{*}$. As $U_{n}U_{n}^{*}$ is self-adjoint, $\beta_{n,j_{n}}=\lVert U_{n}U_{n}^{*}\rVert$ by Lemma A.3, showing $\beta_{n,j_{n}}\rightarrow\infty$. Let $\varepsilon>0$. By density of $\operatorname{ran}{\mathcal{C}^{1/2}_{\textup{pos}}}$, for each $j\leq m$ we can choose $k_{j}\in\mathcal{H}$ which satisfies $\lVert\psi_{1,j}-\mathcal{C}_{\textup{pos}}^{1/2}k_{j}\rVert\leq\varepsilon$. Let us decompose $\psi_{n,j_{n}}=\sum_{j=1}^{m}\langle\psi_{n,j_{n}},\psi_{1,j}\rangle\psi_{1,j}$ and define $h_{n}\coloneqq\sum_{j=1}^{m}\langle\psi_{n,j_{n}},\psi_{1,j}\rangle k_{j}$. Note that $\lVert h_{n}\rVert\leq C$ with $C\coloneqq m\max_{j}\lVert k_{j}\rVert$ by the Cauchy–Schwarz inequality. By further application of the Cauchy–Schwarz inequality,
@@
By Cauchy–Schwarz, the bound $\lVert h_{n}\rVert\leq C$, the fact that $\mathcal{C}_{\textup{pr}}^{1/2}$ is self-adjoint, the given diagonalisation of $U_{n}U_{n}^{*}$ and the previous lower bound,
@@
Let $f\in\mathscr{F}$. By Proposition 4.18, $J_{f}$ is Gateaux differentiable. It follows from [5, Theorem 12.4.5 (i)] that local minimisers of $J_{f}$ are stationary points, i.e. have Gateaux derivative equal to 0. The idea of the proof is to find among all stationary points of $J_{f}$ a stationary point that minimises $J_{f}$, and use the coercivity of $J_{f}$ over finite dimensional subspaces of $\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ to conclude that this stationary point is a global minimiser.
@@
Hence, if $U$ is a stationary point, then $\gamma_{i}U^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i}=0$ for all $i$. Conversely, if $\gamma_{i}U^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i}=0$ for all $i$, then for each $i$ it holds that $U^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i}=0$ or $\gamma_{i}=0$, showing that $f^{\prime}(\gamma_{i})\langle\mathcal{C}_{\textup{pos}}^{1/2},VU^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i}\rangle=0$ for all $i$ and all $V$. Hence $U$ is stationary if and only if $\gamma_{i}U^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i}=0$ for all $i$.
@@
By Lemma 4.19, $\hat{U}$ minimizes $J_{f}$ over $\mathcal{B}(\mathbb{R}^{r},\mathcal{V})$ for the space $\mathcal{V}$ defined above. Furthermore, if $\tilde{U}\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ with $\operatorname{ran}{\tilde{U}}\not\subset\mathcal{V}$, then $\tilde{U}$ is not a stationary point of $J_{f}$, because a necessary condition for $\tilde{U}$ to be a stationary point of $J_{f}$ is that $\operatorname{ran}{\tilde{U}}\subset\mathcal{V}$. On the other hand, since $\hat{U}\in\mathcal{B}(\mathbb{R}^{r},\mathcal{V})\subset\mathcal{B}(\mathbb{R}^{r},\mathcal{V}+\operatorname{ran}{\tilde{U}})$, since $\tilde{U}\in\mathcal{B}(\mathbb{R}^{r},\mathcal{V}+\operatorname{ran}{\tilde{U}})$ and since $J_{f}$ is coercive over $\mathcal{B}(\mathbb{R}^{r},\mathcal{V}+\operatorname{ran}{\tilde{U}})$ by Lemma 4.19, it follows that $J_{f}(\tilde{U})>J_{f}(\hat{U})$. Thus, $\hat{U}$ is a global minimiser of $J_{f}$. Since $J_{f}(\hat{U})=\mathcal{L}_{f}(\mathcal{C}_{\textup{pos}}\|(\mathcal{C}_{\textup{pr}}^{-1}+\hat{U}\hat{U}^{*})^{-1})$ by (17), it follows that the operator $\mathcal{P}^{\textup{opt}}_{r}$ in (21) solves Problem 4.4.
@@
By Corollary 4.8(ii), $(\mathcal{P}^{\textup{opt}}_{r})^{-1}$ solves Problem 4.3. It remains to show that $\mathcal{C}^{\textup{opt}}_{r}$ defined in (22) satisfies $\mathcal{C}^{\textup{opt}}_{r}=(\mathcal{P}^{\textup{opt}}_{r})^{-1}$. Since taking the inverse is a bijective operation, uniqueness of $\mathcal{C}^{\textup{opt}}_{r}$ is then implied by uniqueness of $\mathcal{P}^{\textup{opt}}_{r}$. Now, by (21), Lemma A.9 with $\delta_{i}\leftarrow\lambda_{i}$, and (22),

### v4

@@ B.2 Proofs of Section 4
Given that $xf^{\prime}(x)>0$ for $x\neq 0$, it follows that $f^{\prime}(x)<0$ for $x<0$ and $f^{\prime}(x)>0$ for $x>0$. This implies, by continuity of $f$, that $f^{\prime}(0)=0$. Hence $f$ has a global minimum only at $x=0$ and $f\geq 0$. Thus, also $\mathcal{L}_{f}\geq 0$. By the Lipschitz assumption on $f^{\prime}$ at 0, there exists $\varepsilon\in(0,1)$ and $M_{0}>0$ such that $f^{\prime}(x)=f^{\prime}(x)-f^{\prime}(0)\leq M_{0}\lvert x\rvert$ for $\lvert x\rvert\leq\varepsilon$. For $\lvert y\rvert\leq\varepsilon$,
@@
$\mathcal{C}^{-1}=\mathcal{C}_{1}^{-1/2}\mathfrak{X}^{-1}\mathcal{C}_{1}^{-1/2}=\mathcal{C}_{1}^{-1/2}\left(I+\sum_{i=1}^{\operatorname{rank}\left({K}\right)}\frac{d_{i}^{2}}{1-d_{i}^{2}}e_{i}\otimes e_{i}\right)\mathcal{C}_{1}^{-1/2}=\mathcal{C}_{1}^{-1/2}(I+ZZ^{*})\mathcal{C}_{1}^{-1/2},$
@@
Let $U\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$. We first show that the linear map $\mathcal{B}(\mathbb{R}^{r},\mathcal{H})\rightarrow L_{2}(\mathcal{H})_{\mathbb{R}}$ given by $V\mapsto\mathcal{C}_{\textup{pos}}^{1/2}(UV^{*}+VU^{*})\mathcal{C}_{\textup{pos}}^{1/2}$ is bounded, and then identify this map as the Fréchet derivative of $g$ at $U$. Let $V\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$. By Lemma A.6, $\dim{\ker{V^{*}}^{\perp}}=\operatorname{rank}\left({V}\right)\leq r$. Then there exists an ONB $(e_{i})_{i}$ of $\mathcal{H}$ for which $\operatorname{span}{\left(e_{i},\ i\leq r\right)}$ contains $\ker{V^{*}}^{\perp}$. We have,
$\lVert UV^{*}\rVert_{L_{2}(\mathcal{H})}^{2}=\sum_{i}\lVert UV^{*}e_{i}\rVert^{2}=\sum_{i=1}^{r}\lVert UV^{*}e_{i}\rVert^{2}\leq\sum_{i=1}^{r}\lVert U\rVert^{2}\lVert V^{*}\rVert^{2}=r\lVert U\rVert^{2}\lVert V\rVert^{2},$
where we use consecutively the definition of the $L_{2}(\mathcal{H})$-norm, the inclusion $\ker{V^{*}}^{\perp}\subset\operatorname{span}{\left(e_{1},\ldots,e_{r}\right)}$, the definition of the operator norm, and $\lVert V^{*}\rVert=\lVert V\rVert$ by [8, Proposition VI.1.4(b)]. This also shows $\lVert VU^{*}\rVert_{L_{2}(\mathcal{H})}\leq\sqrt{r}\lVert V\rVert\lVert U\rVert$ and $\lVert VV^{*}\rVert_{L_{2}(\mathcal{H})}\leq\sqrt{r}\lVert V\rVert^{2}$. Thus, using the triangle inequality and the fact $\lVert TS\rVert_{L_{2}(\mathcal{H})}=\lVert ST\rVert_{L_{2}(\mathcal{H})}\leq\lVert T\rVert\lVert S\rVert_{L_{2}(\mathcal{H})}$ for any $T,S\in L_{2}(\mathcal{H})$,
$\lVert\mathcal{C}_{\textup{pos}}^{1/2}(UV^{*}+VU^{*})\mathcal{C}_{\textup{pos}}^{1/2}\rVert_{L_{2}(\mathcal{H})}\leq\lVert\mathcal{C}_{\textup{pos}}^{1/2}\rVert^{2}\lVert UV^{*}+VU^{*}\rVert_{L_{2}(\mathcal{H})}\leq 2\sqrt{r}\lVert\mathcal{C}_{\textup{pos}}^{1/2}\rVert^{2}\lVert U\rVert\lVert V\rVert.$
It follows that $V\mapsto\mathcal{C}_{\textup{pos}}^{1/2}(UV^{*}+VU^{*})\mathcal{C}_{\textup{pos}}^{1/2}$ is bounded. We have by (19),
$g(U+V)-g(U)=\mathcal{C}_{\textup{pos}}^{1/2}((U+V)(U+V)^{*}-UU^{*})\mathcal{C}_{\textup{pos}}^{1/2}=\mathcal{C}_{\textup{pos}}^{1/2}(VV^{*}+UV^{*}+VU^{*})\mathcal{C}_{\textup{pos}}^{1/2}.$
Using once more the fact $\lVert TS\rVert_{L_{2}(\mathcal{H})}=\lVert ST\rVert_{L_{2}(\mathcal{H})}\leq\lVert T\rVert\lVert S\rVert_{L_{2}(\mathcal{H})}$ for any $T,S\in L_{2}(\mathcal{H})$, and the bound $\lVert VV^{*}\rVert_{L_{2}(\mathcal{H})}\leq\sqrt{r}\lVert V\rVert^{2}$ proven above, it follows that
$\lVert g(U+V)-g(U)-\mathcal{C}_{\textup{pos}}^{1/2}(UV^{*}+VU^{*})\mathcal{C}_{\textup{pos}}^{1/2}\rVert_{L_{2}(\mathcal{H})}$ $=\lVert\mathcal{C}_{\textup{pos}}^{1/2}(VV^{*})\mathcal{C}_{\textup{pos}}^{1/2}\rVert_{L_{2}(\mathcal{H})}$
$\leq\lVert\mathcal{C}_{\textup{pos}}^{1/2}\rVert^{2}\lVert VV^{*}\rVert_{L_{2}(\mathcal{H})}\leq\sqrt{r}\lVert\mathcal{C}_{\textup{pos}}^{1/2}\rVert^{2}\lVert V\rVert^{2}.$
@@
where for the second equation we use (29). By definition of $P_{m}$, $\left(P_{m}F^{\prime}(\Lambda(X))\right)_{i}=F^{\prime}(\Lambda(X))_{i}$ for $i\leq m$. Hence,
@@
Let $U,V\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$. Define
$\mathcal{Z}\coloneqq\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}}+\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}(UV^{*}+VU^{*})\mathcal{C}_{\textup{pos}}^{1/2}}+\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}VV^{*}\mathcal{C}_{\textup{pos}}^{1/2}}+\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}H\mathcal{C}_{\textup{pos}}^{1/2}}$
and $\mathcal{W}\coloneqq\{X\in L_{2}(\mathcal{H})_{\mathbb{R}}:\ \operatorname{ran}{X}\subset\mathcal{Z}\}\subset L_{2}(\mathcal{H})_{\mathbb{R}}$. Then $\dim{\mathcal{Z}}<\infty$ since $U$, $V$ and $H$ are finite-rank, and $\operatorname{ran}{g(U+tV)}\subset\mathcal{Z}$ for all $t\in\mathbb{R}$ by definition (19) of $g$, hence $g(U+tV)\in\mathcal{W}$ for all $t\in\mathbb{R}$. Thus, $F_{f}\circ\Lambda\circ g(U+tV)={\left.\kern-1.2pt(F_{f}\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}\circ g(U+tV)$ for all $t\in\mathbb{R}$. By Lemma 4.11 and Proposition 4.16, ${\left.\kern-1.2pt(F_{f}\circ\Lambda)\vphantom{\big{|}}\right|_{\mathcal{W}}}$ is Fréchet differentiable. By Lemma 4.11, $g$ is Fréchet differentiable. In particular, $g$ is Gateaux differentiable at $U$ in the direction $V$. Hence, by the chain rule, $J_{f}$ is Gateaux differentiable at $U$ in the direction $V$. To compute the derivative, we recall that $(e_{j}\otimes e_{k})_{j,k}$ is an ONB of $L_{2}(\mathcal{H})$. The Gateaux derivative of $J_{f}$ at $U$ in the direction $V$ is
@@
Let us write $m\coloneqq\dim{\mathcal{V}}$. For each $n$, let us diagonalise $U_{n}U_{n}^{*}=\sum_{j=1}^{m}\beta_{n,j}\psi_{n,j}\otimes\psi_{n,j}$, where $(\psi_{n,j})_{j=1}^{m}$ forms an ONB of $\mathcal{V}$ and $\beta_{n,j}\geq 0$. Define $j_{n}\coloneqq\operatorname*{arg\,max}_{j\leq m}\beta_{n,j}$, so that $\beta_{n,j_{n}}$ is the largest eigenvalue of $U_{n}U_{n}^{*}$. As $U_{n}U_{n}^{*}$ is self-adjoint, $\beta_{n,j_{n}}=\lVert U_{n}U_{n}^{*}\rVert$ by Lemma A.3, showing $\beta_{n,j_{n}}\rightarrow\infty$. Let $\varepsilon>0$. By density of $\operatorname{ran}{\mathcal{C}^{1/2}_{\textup{pos}}}$, for each $j\leq m$ we can choose $k_{j}\in\mathcal{H}$ which satisfies $\lVert\psi_{1,j}-\mathcal{C}_{\textup{pos}}^{1/2}k_{j}\rVert\leq\varepsilon$. Let us decompose $\psi_{n,j_{n}}=\sum_{j=1}^{m}\langle\psi_{n,j_{n}},\psi_{1,j}\rangle\psi_{1,j}$ and define $h_{n}\coloneqq\sum_{j=1}^{m}\langle\psi_{n,j_{n}},\psi_{1,j}\rangle k_{j}$. Note that $\lVert h_{n}\rVert\leq C$ with $C\coloneqq m\max_{j}\lVert k_{j}\rVert$ by the Cauchy–Schwarz inequality. By further application of the Cauchy–Schwarz inequality,
@@
By the Cauchy–Schwarz inequality, the bound $\lVert h_{n}\rVert\leq C$, the fact that $\mathcal{C}_{\textup{pr}}^{1/2}$ is self-adjoint, the given diagonalisation of $U_{n}U_{n}^{*}$ and the previous lower bound,
@@
Let $f\in\mathscr{F}$. By Proposition 4.18, $J_{f}$ is Gateaux differentiable. It follows from [5, Theorem 12.4.5 (i)] that local minimisers of $J_{f}$ are stationary points, i.e. have Gateaux derivative equal to 0. The idea of the proof is to find among all stationary points of $J_{f}$ the stationary points that minimise $J_{f}$, and use the coercivity of $J_{f}$ over finite dimensional subspaces of $\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ to conclude that these stationary points are global minimisers. We then relate these minimisers to the solutions of Problems 4.4 and 4.3.
Step 1: characterisation of the stationary points of $J_{f}$.
@@
Hence, if $U$ is a stationary point, then $\gamma_{i}U^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i}=0$ for all $i$. Conversely, if $\gamma_{i}U^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i}=0$ for all $i$, then for each $i$ it holds that $U^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i}=0$ or $\gamma_{i}=0$, showing that $f^{\prime}(\gamma_{i})\langle\mathcal{C}_{\textup{pos}}^{1/2}e_{i},VU^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i}\rangle=0$ for all $i$ and all $V$. Hence $U$ is a stationary point of $J_{f}$ if and only if $\gamma_{i}U^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i}=0$ for all $i$.
@@
Step 2: computation of the stationary points of $J_{f}$ with minimal value of $J_{f}$.
@@
Step 3: identification of the stationary points with minimal value of $J_{f}$ as the minimisers of $J_{f}$.
To prove that the minima of $J_{f}$ are precisely the stationary points $\hat{U}$ defined in Step 2, we first show for fixed $U\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ that ${\left.\kern-1.2ptJ_{f}\vphantom{\big{|}}\right|_{\mathcal{B}(\mathbb{R}^{r},\mathcal{V}+\operatorname{ran}{U})}}^{\prime}(U)=0$ implies $J_{f}^{\prime}(U)=0$. Using Lemma 4.10, we can diagonalise $g(U)=\sum_{i}\gamma_{i}e_{i}\otimes e_{i}$ with $(e_{i})_{i}\subset\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}}$. Using the fact that $\mathcal{C}_{\textup{pos}}^{1/2}H\mathcal{C}_{\textup{pos}}^{1/2}=\sum_{i}(-\lambda_{i})v_{i}\otimes v_{i}$, and using the definition of $\mathcal{V}$, it follows that $\mathcal{C}_{\textup{pos}}^{1/2}\mathcal{V}=\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}H\mathcal{C}_{\textup{pos}}^{1/2}}$, and by Lemma A.8, $\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}}=\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}U}$. Thus, for each $j$ for which $\gamma_{j}\not=0$, the identity $\gamma_{j}e_{j}=g(U)e_{j}=\mathcal{C}_{\textup{pos}}^{1/2}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{j}-\mathcal{C}_{\textup{pos}}^{1/2}H\mathcal{C}_{\textup{pos}}^{1/2}e_{j}$ implies that $e_{j}\in\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}UU^{*}\mathcal{C}_{\textup{pos}}^{1/2}}+\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}H\mathcal{C}_{\textup{pos}}^{1/2}}\subset\operatorname{ran}{\mathcal{C}_{\textup{pos}}^{1/2}U}+\mathcal{C}_{\textup{pos}}^{1/2}\mathcal{V}$. Hence $\mathcal{C}_{\textup{pos}}^{-1/2}e_{j}\in\operatorname{ran}{U}+\mathcal{V}$. Now, by the expression of the derivative of $J_{f}$ in Proposition 4.18, ${\left.\kern-1.2ptJ_{f}\vphantom{\big{|}}\right|_{\mathcal{B}(\mathbb{R}^{r},\mathcal{V}+\operatorname{ran}{U})}}^{\prime}(U)=0$ implies $02\sum_{i}f^{\prime}(\gamma_{i})\langle\mathcal{C}_{\textup{pos}}^{1/2}e_{i},VU^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i}\rangle=$ 0 for all $V\in\mathcal{B}(\mathbb{R}^{r},\mathcal{V}+\operatorname{ran}{U})$. For any $\varphi\in\mathbb{R}^{r}$, $V\coloneqq\mathcal{C}_{\textup{pos}}^{-1/2}e_{j}\otimes\varphi\in\mathcal{B}(\mathbb{R}^{r},\mathcal{V}+\operatorname{ran}{U})$ and hence $0=2\sum_{i}f^{\prime}(\gamma_{i})\langle\mathcal{C}_{\textup{pos}}^{1/2}e_{i},\mathcal{C}_{\textup{pos}}^{-1/2}e_{j}\rangle\langle\varphi,U^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i}\rangle=2f^{\prime}(\gamma_{j})\langle\varphi,U^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{j}\rangle$. Since $\gamma_{j}\not=0$, $f^{\prime}(\gamma_{j})\not=0$ by Lemma 4.1(i). Thus, $U^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{j}=0$. We conclude that $\gamma_{i}U^{*}\mathcal{C}_{\textup{pos}}^{1/2}e_{i}=0$ for all $i$. As was shown in Step 1 of the proof, it then holds that $J_{f}^{\prime}(U)$.
@@
By Lemma 4.19, $\hat{U}$ minimises $J_{f}$ over $\mathcal{B}(\mathbb{R}^{r},\mathcal{V})$ for the space $\mathcal{V}$ defined above. Furthermore, if $\tilde{U}\in\mathcal{B}(\mathbb{R}^{r},\mathcal{H})$ with $\operatorname{ran}{\tilde{U}}\not\subset\mathcal{V}$, then $\tilde{U}$ is not a stationary point of $J_{f}$, because a necessary condition for $\tilde{U}$ to be a stationary point of $J_{f}$ is that $\operatorname{ran}{\tilde{U}}\subset\mathcal{V}$, by Step 1. By the previous paragraph, $\tilde{U}$ is not a stationary point of $J_{f}$ restricted to $\mathcal{B}(\mathbb{R}^{r},\mathcal{V}+\operatorname{ran}{\tilde{U}})$. Since $\hat{U}\in\mathcal{B}(\mathbb{R}^{r},\mathcal{V})\subset\mathcal{B}(\mathbb{R}^{r},\mathcal{V}+\operatorname{ran}{\tilde{U}})$, since $\tilde{U}\in\mathcal{B}(\mathbb{R}^{r},\mathcal{V}+\operatorname{ran}{\tilde{U}})$ and since $J_{f}$ is coercive over $\mathcal{B}(\mathbb{R}^{r},\mathcal{V}+\operatorname{ran}{\tilde{U}})$ by Lemma 4.19, it follows that $J_{f}(\tilde{U})>J_{f}(\hat{U})$. Thus, $\hat{U}$ is a global minimiser of $J_{f}$.
Step 4: identification of the solutions of Problems 4.4 and 4.3.
@@
Since $J_{f}(\hat{U})=\mathcal{L}_{f}(\mathcal{C}_{\textup{pos}}\|(\mathcal{C}_{\textup{pr}}^{-1}+\hat{U}\hat{U}^{*})^{-1})$ by (17), it follows that the operator $\mathcal{P}^{\textup{opt}}_{r}$ in (21) solves Problem 4.4. By Corollary 4.8(ii), $(\mathcal{P}^{\textup{opt}}_{r})^{-1}$ solves Problem 4.3. It remains to show that $\mathcal{C}^{\textup{opt}}_{r}$ defined in (22) satisfies $\mathcal{C}^{\textup{opt}}_{r}=(\mathcal{P}^{\textup{opt}}_{r})^{-1}$. Since taking the inverse is a bijective operation, uniqueness of $\mathcal{C}^{\textup{opt}}_{r}$ is then implied by uniqueness of $\mathcal{P}^{\textup{opt}}_{r}$. Now, by (21), Lemma A.9 with $\delta_{i}\leftarrow\lambda_{i}$, and (22),